OpenStudy (anonymous):

1. –(10)–1 (1 point) a. -1/10 b.1/-1^10 c.1/10 d.10

OpenStudy (anonymous):

ITS D 10

OpenStudy (anonymous):

negative times a negative equals a positive

OpenStudy (anonymous):

i figured but got so confuse

OpenStudy (anonymous):

4. Is the number written in scientific notation? If not, explain.5.1 × 10–8 (1 point) Yes; the number is written in scientific notation. No; the first factor is not a number between 1 and 10. No; it is not written as a number times a power of 10. 5. What is the number written in standard notation? 5.71 × 10–3 (1 point) –171.3 0.0571 0.000571 0.00571 Try not to give me answer help lol

OpenStudy (anonymous):

For number 4 look at 5.1 is that in scientific notation?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

Why not?

OpenStudy (anonymous):

its not powered nor do it have zeros

OpenStudy (anonymous):

Exactly so the answer to 4 is what? A B or C

OpenStudy (anonymous):

C :)))

OpenStudy (anonymous):

Good Job, Now look at number 5.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

How do you know whether it is in standard notation or not?

OpenStudy (anonymous):

if its written out

OpenStudy (anonymous):

without being powered

OpenStudy (anonymous):

So... as an example: Typically, it is just the number written out. So for the number 725: Scientific Notation = 7.25 x 102 Expanded Notation = 700 + 20 + 5 Standard Notation = 725

OpenStudy (anonymous):

So, what is the standard notation of 5.71 × 10–3

OpenStudy (anonymous):

hold on lol

OpenStudy (anonymous):

If you multiply the two on a calculator what does it give you?

OpenStudy (anonymous):

d

OpenStudy (anonymous):

the answer is D

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

Anything else?

OpenStudy (anonymous):

I am just trying to catch up and needed help so i am going to stay on for a while until i have to go to church

OpenStudy (anonymous):

Okay, If you need anymore help fan me and you will be able to get ahold of me or just reply to this one and i will come help you. Are you new?

OpenStudy (anonymous):

Yes i am new lol

OpenStudy (anonymous):

some more problems: 7. What is the expression written using each base only once? (–4)–6 · (–4)7 (1 point) 413 –4 13 1

OpenStudy (anonymous):

Okay, Do you know how to get around or need any help. Im on the list of becoming an Ambassador on here so I can help with a lot.

OpenStudy (anonymous):

no as long as i dont cheat and get proper help ;/ i'll be fine lol

OpenStudy (anonymous):

I dont get number 7 please rephrase it

OpenStudy (anonymous):

7. What is the expression written using each base only once? (–4)–6 X (–4)7 (1 point) 413 –4 13 1

OpenStudy (anonymous):

Those are whole numbers not expressions though??

OpenStudy (anonymous):

i got the answer to it

OpenStudy (anonymous):

8. Find the simplified form of the expression. Give your answer in scientific notation. (3 × 106)(8 × 10–4) (1 point) 1.1 × 103 2.4 × 103 1.1 × 10–23 2.4 × 10–23

OpenStudy (anonymous):

So... as an example: Typically, it is just the number written out. So for the number 725: Scientific Notation = 7.25 x 102 Expanded Notation = 700 + 20 + 5 Standard Notation = 725

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Use PEMDAS to find your answer

OpenStudy (anonymous):

Are you sure its not 107 in the first parenthesis?

OpenStudy (anonymous):

Its 4x103dont know how to help you find the answer though... that looks funny

OpenStudy (anonymous):

i just submitted it

OpenStudy (anonymous):

n^6/n^2 9 (1 point) a. 1/n^4 b.n^4 c.n^8 d.n12

OpenStudy (anonymous):

what do you do when you divide powers?

OpenStudy (anonymous):

n^6/n^2 a. 1/n^4 b.n^4 c.n^8 d.n12

OpenStudy (precal):